# and application to tori over local fields that split over cyclic extensions

### v2

@@ 1. Introduction
Let $G$ be a finite group and AA be a $G$-module. We consider the homogeneous cochain complex $F^{\scriptscriptstyle\bullet}$ for AA (see [3, Appendix E]):
@@
We also consider the inhomogeneous cochain complex $C^{\scriptscriptstyle\bullet}$ for AA (see [3, Appendix E]):
@@
see [4, Section IV.7] or Section 2 below. In Section 2, using the canonical isomorphism between the homogeneous and the inhomogeneous complexes, we give formulas for the cup product pairing in terms of inhomogeneous cochains
@@ 1.2. Tori over local fields
(3) $H^{-1}(G,X){\ \xrightarrow{\makebox[11.99998pt]{\scriptsize${\raisebox{-3.01389pt}{$\sim$}}$}}\ }H^{1}(G,X\otimes L^{\times}).$
@@
denote the fundamental class; see, for instance, [5, Definition 9.1]. The isomorphism (3) is given by cup product with $u_{L/K}$.
@@
$e\mapsto(e,L/K)\colon\ K^{\times}\to K^{\times}/{\rm Nm}(L^{\times}){\ \xrightarrow{\makebox[12.26656pt]{\scriptsize${\raisebox{-3.01389pt}{$\sim$}}$}}\ }G$
is the local reciprocity homomorphism; see [4, Section VI.2.2]. Then the fundamental class $u_{L/K}\in H^{2}(G,L^{\times})$ is represented by the following cocycle $a\colon G\times G\to L^{\times}$:
@@
The plan of the note is as follows. In Section 2 we write the formulas for cup product in terms of homogeneous and inhomogeneous cocycles. In Section 3 we prove Theorem $1.$ In Section 4 we prove Theorem 1.2. The arXiv version [1] of this note contains Appendix A where we provide details of calculations giving the “inhomogeneous” formulas of Section 2.
@@ 2. Cup product in terms of inhomogeneous cocycles
We consider the homogeneous cochain complex (1) for AA, which we denote by $F^{\scriptscriptstyle\bullet}$. We also consider the inhomogeneous complex (2) for AA, which we denote by $C^{\scriptscriptstyle\bullet}$. For details, see [3, Appendix E].
@@
We give formulas for cup product of cochains, homogeneous and inhomogeneous. The formulas for homogeneous cochains follow immediately from [4, Section IV.7, formulas for $\varphi_{p,q}$ on page 107]. We deduce formulas for inhomogeneous cochains from these formulas using formulas (4)–(7). Details of calculations are given in Appendix A in the arXiv version [1] of this note.
@@ 3. The fundamental class of a cyclic extension
defined as follows (compare Milne [6, page 102, above Proposition III.1.9]). We represent an element of $G\times G$ as $(\sigma^{i},\sigma^{j})$ with $i,j\in{\mathds{Z}}$, $0\leq i,j<n$. Then
@@
${\rm inv}_{L/K}\colon H^{2}(G,L^{\times}){\ \xrightarrow{\makebox[11.99998pt]{\scriptsize${\raisebox{-3.01389pt}{$\sim$}}$}}\ }\frac{1}{n}{\mathds{Z}}/{\mathds{Z}}.$
@@
see Harari [5, Proposition 9.3]. By construction, we have $\chi(\sigma)=\frac{1}{n}$. Thus
@@ Appendix A Cup product: calculations
For $p,q\geq 0$, set $m=p$, $n=q$. Then by [4, Section IV.7, page 107] we have
@@
For $p,q\geq 1$, set m=pm=-p, n=qn=-q. Then by [4, Section IV.7, page 107] we have
@@
We take $m=p$, n=pqn=-p-q. Then by [4, Section IV.7, page 107] we have
(12) $f_{-q}(\,g_{1},\ \dots,\ g_{q}\,)=\sum_{s}f^{\prime}_{p}(g_{1},\,s_{1},\ \dots,\ s_{p}\,)\,\otimes\,f^{\prime\prime}_{-p-q}(s_{p},\ \dots,\ s_{1},g_{1},\ \dots,\ g_{q}\,).$
Substituting (5) and (7) in the right-hand side, we obtain
@@
We take m=pqm=-p-q, $n=p$. Then by [4, Section IV.7, page 107] we have
Substituting (7) and (5) in the right-hand side, we obtain
@@
We take $m=p+q$, n=qn=-q. Then by [4, Section IV.7, page 107] we have
(16) $f_{p}(g_{0},\dots,g_{p})=\sum_{s}f^{\prime}_{p+q}(g_{0},\dots,g_{p},s_{1},\dots,s_{q})\,\otimes\,f^{\prime\prime}_{-q}(s_{q},\dots,s_{1}).$
Substituting (5) and (7) in the right-hand side, we obtain
@@
We take m=qm=-q, $n=p+q$. Then by [4, Section IV.7, page 107] we have
(18) $f_{p}(g_{0},\ \dots,\ g_{p})=\sum_{s}f^{\prime}_{-q}(s_{1},\ \dots,\ s_{q})\,\otimes\,f^{\prime\prime}_{p+q}(s_{q},\ \dots,\ s_{1},g_{0},\ \dots,\ g_{p}).$
Substituting (7) and (5) in the right-hand side, we obtain

### v3

@@ 1. Introduction
Let $G$ be a finite group and AA be a $G$-module. We consider the homogeneous cochain complex $F^{\scriptscriptstyle\bullet}$ for AA (see [2, Appendix E]):
@@
We also consider the inhomogeneous cochain complex $C^{\scriptscriptstyle\bullet}$ for AA (see [2, Appendix E]):
@@
see [3, Section IV.7] or Section 2 below. In Section 2, using the canonical isomorphism between the homogeneous and the inhomogeneous complexes, we give formulas for the cup product pairing in terms of inhomogeneous cochains
@@ 1.2. Tori over local fields
(3) $H^{-1}(G,X){\ \xrightarrow{\makebox[9.56667pt]{\scriptsize${\raisebox{-3.01389pt}{$\sim$}}$}}\ }H^{1}(G,X\otimes L^{\times}).$
@@
denote the fundamental class; see, for instance, [4, Definition 9.1]. The isomorphism (3) is given by cup product with $u_{L/K}$.
@@
$e\mapsto(e,L/K)\colon\ K^{\times}\to K^{\times}/{\rm Nm}(L^{\times}){\ \xrightarrow{\makebox[10.13342pt]{\scriptsize${\raisebox{-3.01389pt}{$\sim$}}$}}\ }G$
is the local reciprocity homomorphism; see [3, Section VI.2.2]. Then the fundamental class $u_{L/K}\in H^{2}(G,L^{\times})$ is represented by the following cocycle $a\colon G\times G\to L^{\times}$:
@@
The plan of the note is as follows. In Section 2 we write the formulas for cup product in terms of homogeneous and inhomogeneous cocycles. In Section 3 we prove Theorem $1.$ In Section 4 we prove Theorem 1.2. In Appendix A we provide details of calculations giving the “inhomogeneous” formulas of Section 2.
@@ 2. Cup product in terms of inhomogeneous cocycles
We consider the homogeneous cochain complex (1) for AA, which we denote by $F^{\scriptscriptstyle\bullet}$. We also consider the inhomogeneous complex (2) for AA, which we denote by $C^{\scriptscriptstyle\bullet}$. For details, see [2, Appendix E].
@@
We give formulas for cup product of cochains, homogeneous and inhomogeneous. The formulas for homogeneous cochains follow immediately from [3, Section IV.7, formulas for $\varphi_{p,q}$ on page 107]. We deduce formulas for inhomogeneous cochains from these formulas using formulas (4)–(7). Details of calculations are given in Appendix A.
@@ 3. The fundamental class of a cyclic extension
defined as follows (compare Milne [5, page 102, above Proposition III.1.9]). We represent an element of $G\times G$ as $(\sigma^{i},\sigma^{j})$ with $i,j\in{\mathds{Z}}$, $0\leq i,j<n$. Then
@@
${\rm inv}_{L/K}\colon H^{2}(G,L^{\times}){\ \xrightarrow{\makebox[9.56667pt]{\scriptsize${\raisebox{-3.01389pt}{$\sim$}}$}}\ }\frac{1}{n}{\mathds{Z}}/{\mathds{Z}}.$
@@
see Harari [4, Proposition 9.3]. By construction, we have $\chi(\sigma)=\frac{1}{n}$. Thus
@@ Appendix A Cup product: calculations
For $p,q\geq 0$, set $m=p$, $n=q$. Then by [3, Section IV.7, page 107] we have
@@
For $p,q\geq 1$, set m=pm=-p, n=qn=-q. Then by [3, Section IV.7, page 107] we have
@@
We take $m=p$, n=pqn=-p-q. Then by [3, Section IV.7, page 107] we have
where $s_{i}$ run independently through $G$. Substituting (5) and (7) in the right-hand side, we obtain
@@
We take m=pqm=-p-q, $n=p$. Then by [3, Section IV.7, page 107] we have
(14) $f_{-q}(\,g_{1},\ \dots,\ g_{q}\,)=\sum_{s}f^{\prime}_{-p-q}(g_{1},\ \dots,\ g_{q},s_{1},\ \dots,\ s_{p})\,\otimes\,f^{\prime\prime}_{p}(s_{p},\ \dots,\ s_{1},g_{q})$
where $s_{i}$ run independently through $G$. Substituting (7) and (5) in the right-hand side, we obtain
@@
We take $m=p+q$, n=qn=-q. Then by [3, Section IV.7, page 107] we have
where $s_{i}$ run independently through $G$. Substituting (5) and (7) in the right-hand side, we obtain
@@
We take m=qm=-q, $n=p+q$. Then by [3, Section IV.7, page 107] we have
where $s_{i}$ run independently through $G$. Substituting (7) and (5) in the right-hand side, we obtain